Light travels through an optical fiber (n=1.44) until it reaches the fiber's end and emerges into air. What is the angle of refraction outside the fiber if the angle of incidence on the end of the fiber is 25 degrees?


1
Expert's answer
2022-05-05T13:56:35-0400
"n_i\sin\theta_i=n_r\sin\theta_r"

"\theta_r=\sin^{-1}(\frac{n_i}{n_r}\sin\theta_i)"

"\theta_r=\sin^{-1}(\frac{1.44}{1.00}\sin25^\circ)=37^\circ"